OpenStudy (anonymous):

Please Help me :3 In the below system, solve for y in the first equation. x − 3y = −6 2x − 7y = 10 1/3x + 2 -1/3x + 6 −x + 2 -1/3x + 2

OpenStudy (anonymous):

-3y=-6+x y=(-6/-3)+(x/-3) y=2-(1/3)x or y= -(1/3)x + 2

OpenStudy (anonymous):

Sooo D?

OpenStudy (anonymous):

yap !